OpenStudy (anonymous):

Solve for the variable. -22 = x - 7

OpenStudy (shaik0124):

-22+7=x-7+7 x=-15

OpenStudy (anonymous):

I thought maybe to subtract 22 from 7. But that doesn't seem very...well right...

OpenStudy (cgreenwade2000):

Because you are subtracting 7 from x, You actually want to add it on both sides of the equation.

OpenStudy (anonymous):

think of the equation like a balance... you want to get that " - 7" off the right side, but to keep it balance, you have to add seven to both sides... on the right, -7 + 7 = 0, leaving x. On the left, -22+7 = -15

OpenStudy (anonymous):

This is a miracle but I get it. x)~ LOL. Thanks guys!

OpenStudy (anonymous):



OpenStudy (anonymous):

Don't make that see-saw fall left or right, keep it balanced by doing the same thing to both sides... :) (silly example, but it can keep your mind focused)

OpenStudy (anonymous):

add 7 to the -22 to get -15 and the answer is x= -15